D Question 4 8 pts Consider the following recent financials for XYZ Corporation: Income Statement Balance Sheet Sales7 72,429 Assets 178,884 Debt 44,955 Costs 43,457 Equity 133,929 EBIT 28,972 Taxes@ 11,009 Total 178,884 Total 178,884 38% Net 17,963 Income Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2.577 was ichoc to. nintin a constant navout ratio. Next vears sales are

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,577 was paid, and the company wishes to maintain a constant payout ratio, Next years sales are projected to grow by 24%. What is the pro-forma value for equity? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.). What is the external financing needed using the pro-forma approach? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.). What is the internal growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations). What is the sustainable growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations) Topic: Financial Models (EFN & Growth
Answer #1
Year Recent Year Basis for projections Next year
INCOME STATEMENT:
Sales 72429 +24% 89812
Costs 43457 60% of sales 53887
EBIT 28972 35925
Taxes at 38% 11009 13651
Net income 17963 22273
Dividends (14.35%) 2577 3196
Addition to retained earnings 15386 19077
BALANCE SHEET:
Assets 178884 246.98% of sales 221818
Debt 44955 44955
Equity 133929 +19077 153006
Total debt and equity 178884 197961
EFN 23856
ANSWERS:
Proforma value for equity 153006
EFN required 23856
IGR = ROA*b/(1-ROA*b)
where ROA = return on assets,
b = retention ratio.
ROA = 17963/178884 = 10.04%
Retention ratio = 15386/17963 = 85.65%
IGR = 0.1004*0.8565/(1-0.1004*0.8565) = 9.41%
IGR = ROE*b/(1-ROE*b)
where ROE = return on equity,
b = retention ratio.
ROA = 17963/133929 = 13.41%
Retention ratio = 15386/17963 = 85.65%
IGR = 0.1341*0.8565/(1-0.1341*0.8565) = 12.98%